OpenStudy (anonymous):

Find the slope for this linear equation. y = 6x

OpenStudy (anonymous):

A linear equation looks like y= mx +c where m is the slope. That help?

OpenStudy (anonymous):

not at all, this confuses me so much

OpenStudy (anonymous):

So u compare what you have with the one I gave u. The y is the same in both, right?

OpenStudy (anonymous):

yes?

OpenStudy (anonymous):

so its 6, igot it !

OpenStudy (anonymous):

Just to be sure, what's the c in mine when you look at yours?

OpenStudy (anonymous):

the yinercept? how do u find that

OpenStudy (anonymous):

Well, yes it is the y intercept and you don't have anything in your equation, so what does that mean?

OpenStudy (anonymous):

i dont know :/

OpenStudy (anonymous):

does it =zero?

OpenStudy (anonymous):

y = 6x + 0 y = mx + c

OpenStudy (anonymous):

Yes correct. So what does it mean, then?

OpenStudy (anonymous):

ughh u can solve it? equals 6?

OpenStudy (anonymous):

imean zero

OpenStudy (anonymous):

It means the y intercept is 0, doesn't it? U said that already...

OpenStudy (anonymous):

What that means is that your line passes through the origin 0,0

OpenStudy (anonymous):

Ok thanks

OpenStudy (anonymous):

ur welcome, math-s isn-t hard, u just have to get used to it.....honest.